OpenStudy (anonymous):

Given F1=2xi-2yzj-y^2k and F2=yi-xj, Are these forces conservative?

OpenStudy (anonymous):

I want to use the method of curl F, where the operator multiplied by F would tell me if it is conservative IF it equals 0

OpenStudy (ybarrap):

$$ if~ g(x,y,z)=x^2-y^2z-y^3/3\\ then ~\nabla g=F_1 $$ So F1 is conservative. For F2: $$ \partial M/\partial y=1\\ \partial N/\partial x=-1 $$ Since these are not equal, F2 is not conserviative.

OpenStudy (anonymous):

I need those steps broken down alittle more please if you can

OpenStudy (anonymous):

you should take curl of those force f1 is concerviative because: ((d/dy)(-y^2)-(d/dz)(-2yz))i+((d/dz)(-2x)-(d/dx)(-y^2))j+((d/dx)(-2yz)-(d/dy)(-2x))k -2y+2y+0+0=0 so f1 is concervative

OpenStudy (anonymous):

but f2 is not concervative because curl f2=-2